Jefferson Lab HALL A Logbook for April 03 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 99243 04/30/03 23:57 markowitz shift summary
  • 99241 04/30/03 23:50 LeRose Magnet settings
  • 99237 04/30/03 23:16 LeRose Septum Test
  • 99236 04/30/03 23:01 LeRose Septum test
  • 99222 04/30/03 19:56 LeRose Septum Ramp
  • 99220 04/30/03 19:43 Kelleher target to empty target
  • 99217 04/30/03 19:36 Kelleher Target to single foil
  • 99213 04/30/03 18:38 Markowitz HARP scans
  • 99211 04/30/03 16:23 vasulk Short Carbon foil
  • 99208 04/30/03 16:12 vasulk Lifting Motor lower switch
  • 99206 04/30/03 16:05 Ole Hansen Day Shift Summary
  • 99205 04/30/03 15:57 choi Left arm pion rejector L2
  • 99204 04/30/03 15:11 Jack Gas sensors disabled
  • 99203 04/30/03 15:05 Feuerbach IMPT: NEW SAFETY for entering Detector Huts and Gas Shed
  • 99202 04/30/03 14:57 jpchen UPS checked
  • 99201 04/30/03 14:22 choi Q1 dipole reset etc
  • 99192 04/30/03 13:56 R. Michaels Final flush of ring buffer implemented
  • 99167 04/30/03 12:51 vasulk NMR: 36.8%; EPR 36.1%
  • 99160 04/30/03 09:33 vasulk NMR correction
  • 99159 04/30/03 09:00 brindza septum max current
  • 99158 04/30/03 07:20 Solvignon Target: NMR= 34.1%, EPR= 33.6%
  • 99157 04/30/03 02:56 Solvignon Beam update: not before noon
  • 99156 04/30/03 01:56 Solvignon Target: NMR=30.8%, EPR=30.4%
  • 99154 04/30/03 00:39 A. Deur Work done on target this afternoon
  • 99152 04/30/03 00:08 A. Deur Tungsten piece removed from the Hall
  • Tuesday

  • 99151 04/29/03 22:59 Aidan Kelleh Target RTD 5 alarm
  • 99150 04/29/03 19:20 Aidan Kelleh NMR = 31.06%
  • 99149 04/29/03 18:28 vasulk Septum Magnet Instructions
  • 99147 04/29/03 13:45 choi Left arm pion rejector L2
  • 99145 04/29/03 12:53 R. Michaels L-arm DAQ working again
  • 99129 04/29/03 11:43 vasulk Polarization: 21 %
  • 99128 04/29/03 11:36 choi left arm daq failure
  • 99127 04/29/03 11:35 choi left arm pion rejector L2
  • 99123 04/29/03 10:13 vasulk Spectrum Analyzer
  • 99122 04/29/03 10:10 choi Left arm pion rejector level 2
  • 99121 04/29/03 10:09 vasulk Laser # 5
  • 99120 04/29/03 10:05 vasulk Long Carbon Foil Replaced
  • 99118 04/29/03 07:50 Solvignon Target test summary
  • Monday

  • 99117 04/28/03 19:27 vasulk Target work
  • 99115 04/28/03 15:50 shabetai Day Shift Summary
  • 99114 04/28/03 15:38 Feuerbach Sieve spacer mistakes
  • 99113 04/28/03 08:20 vasulk Gas Sensor Box
  • 99112 04/28/03 08:10 vasulk Owl Shift Summary Part 3
  • 99106 04/28/03 07:31 jpchen background study partial summary
  • 99101 04/28/03 07:11 A. Deur holes !
  • 99086 04/28/03 05:22 vasulk Owl Shift Summary Part 2
  • 99083 04/28/03 05:15 vasulk Correction- Thick sieve slit holes not blocked
  • 99060 04/28/03 04:00 vasulk Background not from Helium-4
  • 99057 04/28/03 03:50 vasulk Controlled Access
  • 99051 04/28/03 02:41 LeRose No Holes!
  • 99050 04/28/03 01:39 vasulk Owl Shift Summary Part 1
  • 99043 04/28/03 01:18 A. Camsonne Target moved to single foil C
  • 99029 04/28/03 00:19 LeRose Magnet Settings
  • 99027 04/28/03 00:03 slifer swing shift summary
  • Sunday

  • 99026 04/27/03 23:32 slifer Septum current scan: HRS Quads off.
  • 99015 04/27/03 22:47 LeRose, Mark No quads
  • 98989 04/27/03 22:10 adaq halog entries
  • 98982 04/27/03 21:42 Adaq halog test (ignore)
  • 98967 04/27/03 21:11 Adaq halog test (ignore)
  • 98955 04/27/03 20:47 Adaq halog test
  • 98944 04/27/03 20:13 Adaq .
  • 98931 04/27/03 19:30 Aidan Kelleh target to long foil position
  • 98928 04/27/03 19:24 Feuerbach Single 1-C Septum scan
  • 98925 04/27/03 19:05 Feuerbach New test - R-HRS quads off
  • 98924 04/27/03 18:57 slifer Septum current scan (no target)
  • 98919 04/27/03 18:38 slifer Experiment status update
  • 98916 04/27/03 18:28 slifer Septum current scan
  • 98903 04/27/03 17:53 slifer Septum current scan
  • 98902 04/27/03 17:50 Aidan Kelleh target moved to empty position
  • 98897 04/27/03 16:55 Aidan Kelleh Q2 trip
  • 98887 04/27/03 16:12 j, singh septum strip chart
  • 98885 04/27/03 16:11 vasulk Shift Summary Part 3
  • 98849 04/27/03 15:22 j. singh BeO removed & collimators added
  • 98846 04/27/03 14:16 j. singh plot of T1 rate vs. septum read current
  • 98843 04/27/03 13:12 Feuerbach Fast sweep of septum current
  • 98832 04/27/03 12:10 Higinbotham 4He at 1atm
  • 98831 04/27/03 12:00 Higinbotham Scaler Data from Beamline Changes
  • 98824 04/27/03 11:26 Higinbotham 1k Rate with Septum Off
  • 98823 04/27/03 11:24 Higinbotham Changing Septum Field
  • 98822 04/27/03 10:59 vasulk Laser # 4 Temperature
  • 98821 04/27/03 10:39 vasulk Shift Summary Part 2
  • 98796 04/27/03 09:48 Higinbotham Target Photo
  • 98795 04/27/03 09:12 vasulk Day Shift Summary Part1
  • 98794 04/27/03 09:07 Higinbotham New Background Spectrum
  • 98793 04/27/03 07:15 A. Camsonne HV channel for EP
  • 98792 04/27/03 07:12 A. Camsonne Septum temperature chart
  • 98790 04/27/03 06:32 R. Michaels owl shift summary
  • 98788 04/27/03 06:22 R. Michaels ask for access
  • 98786 04/27/03 06:19 R. Michaels sieve in, collim in runs
  • 98755 04/27/03 05:24 R. Michaels R-Q2 trip
  • 98754 04/27/03 04:10 R. Michaels sieve slit removed, more BGR study
  • 98734 04/27/03 02:41 R. Michaels various runs to explore BGR.
  • 98726 04/27/03 02:28 A. Camsonne Target : laser 4 weak
  • 98718 04/27/03 02:17 A. Camsonne Target moved to BeO/Single Foil C
  • 98711 04/27/03 02:09 R. Michaels empty target run (*300)
  • 98708 04/27/03 01:53 R. Michaels L-arm momentum change
  • 98707 04/27/03 01:51 Nilanga Optics data and background
  • 98706 04/27/03 01:45 R. Michaels adjustments (no septum -> 6 degr !)
  • 98704 04/27/03 01:27 A. Camsonne Target moved to empty
  • 98699 04/27/03 00:54 A. Camsonne Septum temperature chart
  • 98698 04/27/03 00:45 j. singh swing shift summary
  • 98697 04/27/03 00:41 R. Michaels carbon in, rate did not change
  • 98694 04/27/03 00:32 R. Michaels beam quality checks
  • 98689 04/27/03 00:19 j. singh turned on correction coils
  • 98685 04/27/03 00:14 j. singh right Q2 tripped again
  • Saturday

  • 98685 04/26/03 23:49 j. singh right Q2 tripped again
  • 98684 04/26/03 23:38 j. singh septum ramping up to set 98A
  • 98683 04/26/03 23:09 j. singh beam back
  • 98682 04/26/03 23:08 j. singh trouble turning on septum
  • 98677 04/26/03 22:08 j. singh Beam Permit + Empty Target
  • 98676 04/26/03 22:05 j. singh laser alignment problem solved
  • 98672 04/26/03 21:44 Solvignon Laser alignment
  • 98670 04/26/03 21:39 j. singh right Q2 tripped earlier
  • 98669 04/26/03 21:27 j. singh target field to 0 deg
  • 98663 04/26/03 19:03 j. singh set left/right HRS momenta
  • 98662 04/26/03 17:47 j. singh Target @ Pickup Coils, 270
  • 98659 04/26/03 06:21 Solvignon Final results for target tests and polarization correction
  • 98658 04/26/03 05:43 Solvignon Target: NMR = 23.7%; EPR = 24.1%
  • 98657 04/26/03 05:41 Solvignon Summary of target tests
  • 98656 04/26/03 05:33 Solvignon EPR: counter gate
  • 98632 04/26/03 02:11 Solvignon EPR: good spectrum
  • Friday

  • 98577 04/25/03 17:36 slifer problems with hac solved
  • 98576 04/25/03 17:15 slifer problems with hac
  • 98572 04/25/03 15:56 A. Camsonne Day shift summary
  • 98571 04/25/03 15:40 A. Camsonne Septum temperature chart
  • 98570 04/25/03 15:38 A. Camsonne IOC network connections problem, EPICS frozen
  • 98569 04/25/03 14:53 A. Camsonne Septum IOC rebooted by itself
  • 98568 04/25/03 13:13 A. Camsonne Septum temperature chart : magnet powered
  • 98567 04/25/03 11:58 A. Camsonne Septum IOC back
  • 98566 04/25/03 11:37 A. Camsonne Septum IOC frozen again
  • 98565 04/25/03 11:12 A. Camsonne Target temperatures with lasers on
  • 98564 04/25/03 10:20 Adaq Target : Laser on
  • 98563 04/25/03 09:35 A. Camsonne Target temperature up
  • 98562 04/25/03 09:02 A. Camsonne Septum temperature chart
  • 98561 04/25/03 08:40 A. Camsonne Lasers off
  • 98559 04/25/03 07:57 P. Solvignon Shift summary
  • 98558 04/25/03 06:55 Solvignon Target temperature in empty position
  • 98557 04/25/03 06:30 Solvignon RTD#2 instable readback
  • 98556 04/25/03 04:57 Solvignon Target temperature in pick-up coils position
  • 98555 04/25/03 00:32 Solvignon Septum turned off
  • 98554 04/25/03 00:06 slifer swing shift summary
  • Thursday

  • 98551 04/24/03 23:29 slifer CHL crash
  • 98548 04/24/03 22:50 Reitz ep target not existent
  • 98547 04/24/03 22:01 Chang Rate vs beam position study
  • 98546 04/24/03 21:50 vasulk Adjust Correction Coil Current
  • 98545 04/24/03 21:41 vasulk Target Correction Coils
  • 98544 04/24/03 21:27 A. Deur sieve slit removed, position test
  • 98543 04/24/03 20:42 slifer target lasers off
  • 98532 04/24/03 19:01 Chang Possible target dependent rate
  • 98529 04/24/03 18:48 slifer target to multiple carbon foil
  • 98524 04/24/03 18:26 Feuerbach Single and multi-foil difference
  • 98523 04/24/03 18:08 slifer target to single foil
  • 98522 04/24/03 18:05 slifer run synch. 1288/20288
  • 98521 04/24/03 17:19 slifer fast feedback off
  • 98520 04/24/03 17:18 A. Deur control access
  • 98517 04/24/03 17:03 slifer DAQ prob. run numbers
  • 98516 04/24/03 16:59 A. Deur He3 elastic rate
  • 98515 04/24/03 16:58 slifer left arm DAQ problem
  • 98508 04/24/03 16:45 A. Deur Sieve slit photos
  • 98505 04/24/03 16:32 slifer Hall A tools
  • 98504 04/24/03 16:26 A. Deur beam profile
  • 98503 04/24/03 16:20 slifer target to pol he-3 position
  • 98493 04/24/03 16:07 Saha, A Vach Shift Summary
  • 98492 04/24/03 16:06 A. Saha BTA
  • 98484 04/24/03 15:51 X. Zheng target moved to EMPTY position
  • 98471 04/24/03 15:29 Higinbotham Hysteresis ?
  • 98470 04/24/03 15:26 Higinbotham Sieve Spacing
  • 98469 04/24/03 15:16 vasulk NMR measurement
  • 98468 04/24/03 14:43 vasulk Increase Correction coil current
  • 98467 04/24/03 14:43 A.Vacheret Q2R resetted
  • 98466 04/24/03 14:40 Reitz Bulls Eye Scan: Harp results
  • 98465 04/24/03 14:24 A. Vacheret RQ2 tripped
  • 98464 04/24/03 13:46 vasulk BeO target centered in camera
  • 98462 04/24/03 13:29 Higinbotham One nice delta peak
  • 98461 04/24/03 13:21 Higinbotham How it was tuned
  • 98460 04/24/03 12:48 Gavalya sieve slit inspection
  • 98459 04/24/03 12:37 A. Deur Access, lasers & target collimators
  • 98458 04/24/03 12:35 vasulk Target Controlled Access Activities
  • 98457 04/24/03 12:32 Higinbotham Hall Checks
  • 98456 04/24/03 12:06 X. Zheng target: lasers were turned off at 11am
  • 98455 04/24/03 12:03 X. Zheng target moved to EMPTY position
  • 98452 04/24/03 10:57 X. Zheng target moved to BeO+12C Long Foil position
  • 98451 04/24/03 10:49 A. Vacheret Septum temperature rising
  • 98445 04/24/03 10:20 A.Vacheret Septum temperature rising
  • 98444 04/24/03 10:20 X. Zheng target moved to Pol3He position
  • 98443 04/24/03 10:13 Higinbotham Different Cuts
  • 98442 04/24/03 10:02 Higinbotham Cut on large holes dp looks better
  • 98433 04/24/03 09:55 Higinbotham Sieve Data (no cuts)
  • 98432 04/24/03 09:51 Higinbotham Spectrum with sieve in wrong location
  • 98430 04/24/03 09:49 X. Zheng target moved to reference cell position
  • 98428 04/24/03 09:42 A. Deur left arm data
  • 98424 04/24/03 09:37 A Vacheret raster size check
  • 98420 04/24/03 09:31 X. Zheng possible reason for EMPTY target problem
  • 98417 04/24/03 09:25 X. Zheng pulsed beam
  • 98416 04/24/03 09:19 A. Vacheret resynchronisation of run numbers
  • 98415 04/24/03 09:09 X. Zheng Sth. wrong with target EMPTY position, moved to 12C foils.
  • 98414 04/24/03 09:08 A. Deur Septum temperature strip chart
  • 98413 04/24/03 08:45 Higinbotham Problem with Hall A Quads
  • 98412 04/24/03 08:44 A. Deur raster
  • 98411 04/24/03 08:40 X. Zheng target moved to EMPTY position
  • 98410 04/24/03 08:40 A. Deur Q2 quench
  • 98409 04/24/03 08:39 A. Deur OTR
  • 98408 04/24/03 08:37 A. Deur septum warming up
  • 98406 04/24/03 08:06 Solvignon Shift checklist
  • 98400 04/24/03 07:36 A. Camsonne Target : moved to reference cell
  • 98393 04/24/03 06:41 A. Glamazdin N2 level in left Q3
  • 98392 04/24/03 06:26 A. Camsonne Target : change target gradient coil with septum current
  • 98391 04/24/03 06:16 A. Camsonne NMR measurement
  • 98377 04/24/03 03:38 A. Camsonne Right spectrometer magnets settings
  • 98376 04/24/03 03:35 Eugene Moller results
  • 98375 04/24/03 00:07 Reitz septum temps
  • 98373 04/24/03 00:02 Reitz Shift Summary
  • Wednesday

  • 98372 04/23/03 22:34 Reitz septum quench
  • 98371 04/23/03 22:08 Reitz septum screen
  • 98370 04/23/03 21:58 Reitz septum lost
  • 98369 04/23/03 20:48 slifer turned on gradient coil
  • 98368 04/23/03 20:11 reitz right q2 lost
  • 98316 04/23/03 17:51 vasulk Proteus RTD Positions
  • 98314 04/23/03 17:43 Moffit Parity DAQ: bcm, battery change
  • 98307 04/23/03 17:07 Ole Hansen Shift Summary
  • 98305 04/23/03 16:58 slifer Gradient coil information
  • 98304 04/23/03 16:39 Nanda Compton
  • 98296 04/23/03 13:37 vasulk Controlled Access Activities
  • 98293 04/23/03 12:06 Ole Hansen Day Shift Run Summary part 1
  • 98278 04/23/03 09:57 Ole Hansen Current to 5uA
  • 98273 04/23/03 09:41 Saha Arc Integral Measurement
  • 98268 04/23/03 09:25 vasulk BeO Camera
  • 98266 04/23/03 09:13 vasulk Encoder position for Empty target
  • 98264 04/23/03 09:07 vasulk BeO+12Carbon positions
  • 98260 04/23/03 08:05 Paschke/Solv Shift summary
  • 98242 04/23/03 05:52 paschke IA scan
  • 98239 04/23/03 05:28 paschke parity DAQ checkout
  • 98238 04/23/03 05:25 paschke formula for septum current setpoint
  • 98206 04/23/03 03:53 Solvignon right cerenkov
  • 98200 04/23/03 03:23 Solvignon right cerenkov: PMT#3
  • 98173 04/23/03 02:47 paschke Moved to Empty Target
  • 98153 04/23/03 01:42 paschke Moved to Empty Target
  • 98149 04/23/03 01:31 LeRose Sieve at 6 degrees
  • 98143 04/23/03 00:02 Eugene Run Summary
  • Tuesday

  • 98126 04/22/03 21:47 R. Michaels L-arm filesplitting enabled too
  • 98121 04/22/03 21:30 R. Michaels file splitting enabled
  • 98120 04/22/03 20:59 Eugene Control Access
  • 98096 04/22/03 18:04 Eugene Control access
  • 98090 04/22/03 17:54 kramerk Target to Empty Position
  • 98084 04/22/03 17:24 Moffit Parity: ADC Gain Study at 1 uA
  • 98070 04/22/03 16:59 Eugene Raster Studies
  • 98053 04/22/03 15:52 A. Deur Shift summary
  • 98052 04/22/03 15:46 A. Deur daq faq
  • 98050 04/22/03 15:39 Moffit Parity DAQ: Changed BCM definition
  • 98045 04/22/03 15:30 R. Michaels BTA auto numbers fixed
  • 98043 04/22/03 15:25 LeRose Septum Tests
  • 98042 04/22/03 15:13 A. Deur rate and prescalers
  • 98032 04/22/03 14:32 R. Michaels prl2 missing channels
  • 98029 04/22/03 14:12 A. Deur BTA
  • 98022 04/22/03 14:05 A. Deur beam back
  • 98021 04/22/03 14:04 A. Deur beam down
  • 98020 04/22/03 13:40 A. Deur 10 uA in Hall A
  • 98017 04/22/03 13:11 vasulk Reference cell check
  • 98016 04/22/03 13:04 vasulk Move to Empty Target position
  • 98007 04/22/03 12:13 vasulk Move target to BeO position
  • 98001 04/22/03 11:56 A. Deur OTR in
  • 98000 04/22/03 11:47 A. Deur cw beam
  • 97997 04/22/03 11:10 vasulk Helmholtz Coil Current GUI
  • 97992 04/22/03 10:16 vasulk Target encoder positions
  • 97991 04/22/03 09:59 A. Deur Pulse beam
  • 97986 04/22/03 08:18 A. Camsonne Owl Shift summary
  • 97984 04/22/03 04:31 A. Camsonne EP Ion chamber
  • 97983 04/22/03 04:06 A. Camsonne Pulsed beam for tuning
  • 97982 04/22/03 02:08 Solvignon Longitudinal laser wavelength
  • 97980 04/22/03 00:01 Solvignon RTD readback with 3 lasers at full power
  • 97979 04/22/03 00:01 A. Deur Shift summary
  • Monday

  • 97975 04/21/03 22:38 A. Deur Pulsed beam
  • 97974 04/21/03 22:15 A .Deur Pulse beam
  • 97973 04/21/03 22:04 R. Michaels bta working ...
  • 97970 04/21/03 20:50 A. Deur Hall A sweep is done
  • 97969 04/21/03 20:50 A. Deur Q2 right arm triped
  • 97968 04/21/03 20:37 X. Zheng move target to empty postion
  • 97967 04/21/03 19:37 A. Deur Left arm momentum setting
  • 97966 04/21/03 18:46 X. Jiang left arm pion rejector. most likely a software problem.
  • 97963 04/21/03 18:30 vasulk Target encoder positions
  • 97962 04/21/03 18:10 R. Michaels bta broken
  • 97959 04/21/03 16:03 A. Vacheret shift summary
  • 97905 04/21/03 08:33 Solvignon Detector check
  • 97904 04/21/03 07:58 solvignon Shift summary
  • 97901 04/21/03 07:02 j. singh FIRE? - left shield hut alarm
  • 97895 04/21/03 04:54 j singh mech. switch fixed
  • 97892 04/21/03 04:12 Solvignon temporary target positions
  • 97891 04/21/03 03:16 j singh target FSD problem
  • Sunday

    
    

    Saturday

  • 97884 04/19/03 16:43 R. Michaels Timing adjustments
  • 97878 04/19/03 14:25 atrig Trigger-Download
  • Friday

  • 97859 04/18/03 16:17 R. Michaels plugged in EDTM pulser
  • 97854 04/18/03 16:08 atrig Trigger-Download
  • 97846 04/18/03 14:14 atrig Trigger-Download
  • Thursday

  • 97803 04/17/03 15:48 atrig Trigger-Download
  • Wednesday

  • 97731 04/16/03 16:48 R. Michaels trigger latch (part 2)
  • 0 04/16/03 16:22 no_name no_key
  • 97707 04/16/03 13:21 R. Michaels trigger latch into TDC
  • Tuesday

  • 97689 04/15/03 17:26 R. Michaels plugged in T1 to a TDC
  • 97681 04/15/03 16:41 R. Michaels helicity in ADCs
  • 97678 04/15/03 15:49 atrig Trigger-Download
  • Monday

    
    

    Saturday

    
    

    Thursday

  • 97621 04/10/03 16:01 atrig Trigger-Download
  • Wednesday

  • 97610 04/09/03 16:11 R. Michaels multiple files for run 20585
  • Tuesday

    
    

    Monday

    
    

    Friday

  • 97486 04/04/03 14:42 atrig Trigger-Download
  • 97485 04/04/03 11:55 atrig Trigger-Download
  • 97483 04/04/03 11:50 atrig Trigger-Download
  • 97482 04/04/03 11:48 atrig Trigger-Download
  • 97481 04/04/03 11:48 atrig Trigger-Download